# Detecting Spike Wave Discharges (SWD) using 1-dimensional Residual UNet

### v1

@@ 3. Inverting
where - $x(t)\in\mathbb{R}^{T}$: Original single-channel EEG signal - $\tilde{x}(t)\in\mathbb{R}^{T}$: Augmented EEG signal - $T$: Number of time samples - $\alpha$: Scaling factor - $\sigma^{2}$: Noise variance - $\mathcal{U}(a,b)$: Uniform distribution between $a$ and $b$ - $\mathcal{N}(\mu,\sigma^{2})$: Normal distribution with mean $\mu$ and variance $\sigma^{2}$
We further divided a full day of data (signal and label) into 20 second chunks (2000 time points each), before using them for training our model.
This combination of residual learning, U-Net’s hierarchical temporal representations, and augmentation-based regularization proved essential for achieving robust cross-subject generalization across our 10 test mice at 100Hz temporal resolution.
@@
Dice Loss has become a standard loss function for segmentation tasks across both computer vision and time series applications, making it a natural choice for our seizure detection framework. Originally popularized in medical image segmentation where it effectively addresses class imbalance and spatial overlap measurement (Milletari et al., 2016 [29]), Dice Loss has since been successfully adapted to temporal segmentation tasks including sleep stage classification (Perslev et al., 2019 [17]) and other physiological signal analysis applications. Building on these established successes, we adopted Dice Loss for our cross-subject seizure detection task.
Unlike traditional cross-entropy loss, which treats each pixel or time point independently, Dice Loss considers the global overlap of the segmented regions, making it inherently more robust to class imbalance. This property explains its widespread adoption in both spatial and temporal segmentation tasks. For our application, this meant the model would not simply predict the majority (non-seizure) class but would instead optimize for accurate detection of the minority seizure events.
Mathematically Dice Loss is defined as
$\mathcal{L}_{Dice}=1-\frac{2\sum_{i=1}^{N}p_{i}g_{i}}{\sum_{i=1}^{N}p_{i}+\sum_{i=1}^{N}g_{i}}$
where: $p_{i}$ = predicted probability for pixel $i$, $g_{i}$ = ground truth label for pixel $i$, NN = total number of pixels, $C$ = number of classes, $\epsilon$ = small constant typically $10^{-7}$) for numerical stability, $P$ = predicted segmentation set, $G$ = ground truth segmentation set
@@
Due to the highly imbalanced nature of our data, where seizure signals at the most are 5% of our dataset, we rely on F1-scores, precision and recall to evaluate the performance of our model, as accuracy would not be a reliable metric. Our task involves classifying every time point in the signal as either seizure and non-seizure and the chosen metrics are appropriate for this class imbalanced problem.
@@
where $TP$: True Positives (correctly predicted positive cases) $TN$: True Negatives (correctly predicted negative cases) $FP$: False Positives (incorrectly predicted as positive) $FN$: False Negatives (incorrectly predicted as negative)
@@
To rigorously evaluate the performance of our proposed AugUNet1D architecture, we selected 14 diverse baseline models spanning multiple methodological paradigms and complexity levels. Our selection strategy was designed to provide comprehensive coverage of approaches used in EEG analysis and time series classification tasks. First, we included four traditional machine learning methods, namely Logistic Regression, K-Nearest Neighbors, Decision Tree, and Random Forest, which serve as foundational baselines and represent classical approaches to seizure detection that rely on learned feature representations from the input signals. Second, we incorporated five deep learning architectures of varying complexity: standard CNN and PyramidalCNN for spatial feature extraction, LSTM and bidirectional LSTM (biLSTM) for temporal modeling, and ConvLSTM which combines both spatial and temporal processing. Third, we selected five specialized architectures that have demonstrated success in related physiological signal processing tasks: EEGNet (specifically designed for EEG classification), InceptionTime (a state-of-the-art time series classifier), Xception (an efficient convolutional architecture), vanilla UNet (to isolate the contribution of our residual and augmentation enhancements), and SalientSleepNet (a recent attention-based architecture for sleep stage classification). Finally, we included DETRtime, a transformer-based object detection model adapted for time series, to assess whether modern query-based detection paradigms could be effective for dense temporal segmentation tasks. This diverse baseline selection enables us to systematically evaluate the contributions of different architectural components, including encoder-decoder structures, skip connections, residual learning, recurrent processing, and attention mechanisms, and to determine which design principles are most critical for achieving robust cross-subject seizure detection performance.
@@
Performance scores (F1, recall, precision) of AugUNet1D and Twin Peaks were compared using paired samples t-tests. Differences event features (duration and peak frequency) between manual labeling, AugUNet1D, and Twin Peaks were tested using one-way repeated measures ANOVA with detection method as the main factor. Bonferroni post hoc tests were used to compare groups means between all possible combinations of the three detection methods. All statistical tests were performed in GraphPad Prism 8.0.1 (San Diego, CA). Data are reported as mean ± standard deviation (SD) unless otherwise stated.
@@
All EEG signals underwent identical preprocessing: first, the raw signals were min-max scaled over their amplitude range to normalize signal magnitudes across subjects; second, signals were resampled to 100Hz using torchaudio.transforms.Resample to standardize temporal resolution; and finally, the continuous recordings were segmented into 20-second windows (2000 time points each) to create manageable training examples. The training data from the 8 training mice was randomly partitioned into 95% for model training and 5% for validation, with the validation set used to monitor performance and select optimal model checkpoints during training. The batch size used was 32.
The AugUNet1D model was trained using the Adam optimizer with an initial learning rate of $10^{-3}$ for a maximum of 50 epochs. We employed a CosineAnnealingLinearWarmup learning rate scheduler to facilitate stable training and optimal convergence, which began with a linear warmup phase over the first 500 steps to gradually increase the learning rate from zero to the initial value, preventing early training instability. Following warmup, the scheduler applied cosine annealing over 1000-step cycles with a minimum learning rate of $10^{-5}$ and a decay factor (gamma) of 0.9, allowing the learning rate to oscillate while gradually decreasing over successive cycles to enable fine-tuning of model parameters. To prevent overfitting and reduce unnecessary computational costs, we implemented early stopping with a patience of 10 epochs that monitored the validation Dice loss and terminated training if no improvement was observed for 10 consecutive epochs. This training configuration balanced exploration during early training phases with exploitation during later stages, while the early stopping criterion ensured that models were evaluated at their optimal generalization point rather than after potential overfitting. We report results as mean and standard deviation over 3 runs with random initialization.
@@ Results
As described in model comparison section above, we compare our method against a variety of traditional ML and neural network based methods.
We evaluated our proposed AugUNet1D model against 14 baseline methods and the Twin Peaks approach across 10 held-out test mice, using precision, recall, and F1-score as our primary evaluation metrics. The baseline methods spanned four methodological categories: traditional machine learning algorithms (Logistic Regression, K-Nearest Neighbors, Decision Tree, Random Forest), pure deep learning architectures (CNN, PyramidalCNN, LSTM, biLSTM, ConvLSTM), specialized physiological signal processing networks (EEGNet, InceptionTime, Xception, UNet, SalientSleepNet), and transformer-based detection models (DETRtime). Additionally, we compared against Twin Peaks, the previous state-of-the-art non-machine learning method on this dataset. Table 2 presents the complete performance comparison across all methods and all 10 test mice.
Table 2: Table shows the performance metrics for each model tested across all 10 of our test mice. The ’avg’ columns contains the mean and standard deviation over all mice. We can see that our method performs the best on our dataset.
@@
Table 2 shows our proposed AugUNet1D achieved superior performance with an average F1-score of 0.90 ± 0.01, representing a 29% relative improvement over the Twin Peaks baseline (0.69 F1-score) and a 70% relative improvement over the best general deep learning baseline. Notably, AugUNet1D demonstrated exceptional cross-subject consistency with standard deviations of only 0.01-0.02 across all metrics, the lowest variance among all evaluated methods. The model achieved robust individual performance across all test subjects, with F1-scores ranging from 0.87 to 0.93, indicating successful generalization to held out mice. Furthermore, AugUNet1D maintained a balanced precision-recall trade-off with 0.91 precision and 0.90 recall, compared to Twin Peaks’ 0.67 precision and 0.81 recall, suggesting that our model achieves superior detection accuracy without sacrificing either sensitivity or specificity.
@@
The baseline methods exhibited dramatically different performance profiles that reveal important insights about architectural requirements for cross-subject seizure detection. Traditional machine learning methods performed poorly across the board, with F1-scores ranging from 0.16 for Logistic Regression to 0.33 for Decision Tree. Despite their consistent performance (low variance), these methods fundamentally struggled with the cross-subject temporal segmentation task likely due to their limited capacity to learn complex hierarchical representations from raw EEG signals. The failure of these methods underscores the necessity of deep learning approaches for this challenging problem.
@@
Surprisingly, pure recurrent architectures including LSTM, biLSTM, and ConvLSTM completely failed to learn effective representations, achieving 0.00 F1-scores across all test mice. This catastrophic failure suggests that recurrent networks, despite their theoretical suitability for sequential data, require careful initialization and architectural support to handle cross-subject generalization in high-frequency EEG signals.
Standard convolutional architectures achieved moderate performance, with vanilla CNN reaching 0.47 ± 0.05 F1-score and PyramidalCNN slightly improving to 0.53 ± 0.04 F1-score. However, these models exhibited high variance in precision (0.74 ± 0.21 for CNN), suggesting inconsistent performance across different subjects. The modest improvement of PyramidalCNN over vanilla CNN indicates that hierarchical feature extraction provides some benefit, but still falls far short of what is needed for robust cross-subject detection.
@@
Among specialized architectures, vanilla UNet achieved the highest F1-score of 0.59 ± 0.15, emerging as the strongest general deep learning baseline. However, its high standard deviation of 0.15 reveals significant instability across subjects, highlighting the importance of our residual connection modifications. InceptionTime and Xception both achieved identical performance, demonstrating that multi-scale temporal feature extraction and efficient convolutional designs alone are insufficient to overcome the cross-subject generalization challenge. Most surprisingly, EEGNet achieved only 0.19 ± 0.10 F1-score despite being specifically designed for EEG classification tasks.
SalientSleepNet, despite incorporating attention mechanisms, achieved 0.47 ± 0.36 F1-score with extremely high variance (0.36 standard deviation), indicating that attention alone does not guarantee robust cross-subject generalization.
@@
One recently published method, known as "Twin Peaks", was developed for EEG recorded from the WAG/Rij strain of rats, which have spontaneous SWD, and showed strong performance[23]. To determine how AugUNet1D compares to Twin Peaks at a granular resolution, we applied both methods to our test dataset and compared the performance metrics as well as features of the detected events (i.e. durations and peak frequencies).
@@
In these comparisons, performance was measured on a whole event basis. In other words, if a detected event had any overlap with a manually labeled event, it was considered a true positive. We classified only events that shared no overlap at all as false positives (Figure 4A). Similarly, only manually labeled events that were missed entirely were considered false negatives (Figure 4B). Using this approach AugNet1D had a significantly higher mean F1 score (0.95 ± 0.04) than Twin Peaks (0.71 ± 0.15; t(9) = 5.34, p < 10-3). AugNet1D also showed higher precision (0.96 ± 0.03) than Twin Peaks (0.58 ± 0.18; t(9) = 7.38, p < 10-4). Twin Peaks did have a slightly higher recall (0.96 ± 0.07) than AugUNet1D (0.94 ± 0.06; t(9) = 3.95, p = 0.003), but the magnitude of the effects observed for F1 and precision scores were far larger. Because F1 scores balance precision and recall, we conclude that AugUNet1D outperformed Twin Peaks in our test dataset, having better F1 scores in every recording therein (Figure 4C).
@@
Figure 4: False positives, negatives, confusion matrices and F1 scores for AugUNet1D and Twin Peaks. A) Example of false positive in the dotted blue box. B) Example of false negative in the dotted red box. Blue lines correspond to events detected by AugUNet1D. Red lines correspond to manually labeled events in A and B. C) Confusion matrices and F1 scores for AugUNet1D and Twin Peaks prediction approaches for each mouse in the test dataset. Proportions in colored panels are as follows: Top left, true positives to all positives. Top right, false positives to all positives. Bottom left, false negatives to all negatives. Bottom right, true negatives to all negatives. These values were calculated on a per event basis, meaning that, for example, a predicted SWD was considered true if it had any overlap with an manually labeled SWD and false otherwise.
@@
We also compared features of the events detected by each approach to determine if either method had a bias towards particular types of events. Specifically, we compared the peak frequency of detected events and the average duration of events, which, in our original training data, were similar across time and across mice. A one-way repeated measures ANOVA revealed a significant effect of detection method on event duration (F(2,18) = 92.7, p < 10-4). Post hoc comparison showed that Twin Peaks detected events that were significantly shorter (2.35 ± 0.74s) than Manual (4.88 ± 0.89s; t(9) = 11.7, p < 10-3) and AugUNet1D (4.93 ± 0.82s; t(18) = 11.9, p < 10-3; Figure 5B and 5C). One-way repeated measures ANOVA also showed a significant effect of detection method on peak frequencies (F(2,18) = 18.25, p < 10-4). Twin Peaks detected events with higher (6.58 ± 0.42Hz) frequencies than AugUNet1D (6.13 ± 0.14Hz; t(18) = 5.26, p < 10-3) and Manual (6.14 ± 0.13Hz; t(9) = 5.2, p < 10-3; Figure 5A and 5D). There were no significant difference between AugUNet1D and Manual in duration (t(18) = 0.2, p > 0.99) or peak frequency (t(18) = 0.06, p > 0.99). To summarize, in addition to performing better in terms of overall F1 score, AugUNet1D also detected events with features similar to manually labeled SWDs. Twin Peaks detected events with higher peak frequencies and shorter durations.
@@
Figure 5: Test dataset statistics, spectral features, and comparison across detection methods. A) Peak frequency of events detected manually, with the modified UNet, or using Twin Peaks across all ten mice in the test dataset. B) Same as A but showing event durations. C) Bars show average durations across all ten mice using three different detection methods. Open circles show the median event durations for each mice. D) Same as in C but showing event peak frequencies. The Twin Peaks event detection method shows a bias for shorter events with higher peak frequencies. Error bars show ± standard deviation
@@
We next asked whether events were erroneously detected in specific periods of the recording, such as during noise induced by movement, poor mechanical connection, or cable swinging, or during sleep. Briefly, noise epochs were defined by blocks of time filled with artificially high amplitude EEG signal (see Materials and Methods). These recordings had very few noise epochs overall, making it difficult to rigorously characterize performance during noisy periods, but we observed that Twin Peaks detects more events during noise epochs. This was especially true in a specific recording with many large noise transients (i.e. mouse 2, noise epochs in yellow in Figure 6C). Sleep epochs were identified by increases in the delta (0.1 to 4Hz) power envelope (see Materials and Methods). One-way repeated measures ANOVA revealed a significant effect of detection method on the proportion of events during sleep (F(2,18) = 35.06, p < 10-4). Post hoc comparisons showed that a greater proportion of events during sleep were detected by Twin Peaks (0.46 ± 0.15) than Manual (0.17 ± 0.07; t18 = 7.36 p < 10-4) and AugUNet1D (0.18 ± 0.017; t18 = 7.14 p < 10-4). There was no difference between Manual and AugUNet1D (t(18) = 0.22, p > 0.99, Figure 6D). In conclusion, short non-SWD events are erroneously detected by Twin Peaks during during transient periods of high amplitude noise and during sleep.
Figure 6: Detected events during noise and sleep. A) Example of the Twin Peaks method falsely detecting two events during periods with high-amplitude cable artifact highlighted in yellow. B) Example trace with Twin Peaks falsely detecting many events during sleep which is indicated in blue shading. C) The proportion of events that Twin Peaks detects during sleep is higher than the number of manually detected events and those detected by AugUNet1D. D) Twin Peaks also detects many more events during sleep than manual labelling and AugUNet1D.
@@
AugUNet1D uses several data augmentation strategies during training to diversify the dataset in the hopes of improving generalization to new, unseen data. We conducted an ablation study to assess the individual and combined contributions of each data augmentation strategy: Gaussian noise injection, signal inversion, and amplitude scaling—on the cross-subject seizure detection performance of our residual 1D U-Net architecture. The results reveal critical insights into the role of augmentation in achieving robust generalization. Results are listed in Table 3.
Table 3: Model performance on 400Hz data for different chosen augmentations. We can see scaling provides the most benefit for our use case.
@@
Without any data augmentation, the model achieved high precision (0.9464) but catastrophically low recall (0.2755), resulting in a poor F1-score of 0.4268. This severe imbalance indicates that the unaugmented model learned to be extremely conservative in its predictions, likely defaulting to predicting the majority (non-seizure) class to minimize false positives. While the model could identify seizures with high confidence when it made positive predictions, it missed the vast majority of actual seizure events, rendering it clinically ineffective.
Adding Gaussian noise augmentation alone provided only marginal improvement, increasing recall to 0.3116 and F1-score to 0.4698 while maintaining high precision (0.9544). This modest benefit suggests that noise injection helps the model become slightly more robust to signal variations, but does not fundamentally address the class imbalance problem. Signal inversion augmentation showed slightly greater impact, improving recall to 0.3631 and F1-score to 0.5267 with precision of 0.9591.
In contrast, amplitude scaling augmentation alone produced a dramatic transformation in model behavior, achieving a recall of 0.8442 and F1-score of 0.8609. While precision decreased moderately to 0.8783, the substantial gain in recall demonstrates that scaling augmentation is by far the most critical component for cross-subject generalization.
Combining all three augmentation strategies yielded the best overall performance with an F1-score of 0.8848, achieving an optimal balance between precision (0.9092) and recall (0.8618). The combined approach outperformed even amplitude scaling alone, indicating that while scaling addresses the primary challenge of amplitude variability, the additional regularization from noise injection and the polarity invariance from signal inversion provide complementary benefits that further improve model robustness.

### v2

@@ 3. Inverting
where
- $x(t)\in\mathbb{R}^{T}$: Original single-channel EEG signal
- $\tilde{x}(t)\in\mathbb{R}^{T}$: Augmented EEG signal
- $T$: Number of time samples
- $\alpha$: Scaling factor
- $\sigma^{2}$: Noise variance
- $\mathcal{U}(a,b)$: Uniform distribution between $a$ and $b$
- $\mathcal{N}(\mu,\sigma^{2})$: Normal distribution with mean $\mu$ and variance $\sigma^{2}$
We further divided a full day of data (signal and label) into 20 second epochs (2000 time points each), before using them for training our model.
Ultimately, the combination of residual learning, U-Net’s hierarchical temporal representations, and augmentation-based regularization proved essential for achieving robust cross-subject generalization across our 10 test mice at 100Hz temporal resolution.
@@
Dice Loss is now a standard loss function for segmentation tasks across both computer vision and time series applications, making it a natural choice for our seizure detection framework. Originally popularized in medical image segmentation to effectively address class imbalance and spatial overlap measurement[29]), Dice Loss has since been successfully adapted to temporal segmentation tasks including sleep stage classification[17] and other physiological signal analysis applications. Building on these established successes, we adopted Dice Loss for our cross-subject seizure detection task.
Unlike traditional cross-entropy loss, which treats each pixel or time point independently, Dice Loss considers the global overlap of the segmented regions, making it inherently more robust to class imbalance. This property explains its widespread adoption in both spatial and temporal segmentation tasks. For our application, adopting this approach would not simply predict the majority (non-seizure) class but would instead optimize for accurate detection of the more infrequent event class (i.e. SWDs).
@@
Due to the highly imbalanced nature of our data–where seizures comprise, at most, 5% of our dataset–we rely on F1-scores, precision and recall to evaluate the performance of our model, as accuracy would not serve as a reliable metric. Our task involves classifying every time point in the signal as either seizure or non-seizure, and the chosen metrics are appropriate for this class imbalanced problem.
@@
where
$\mathit{TP}$ $:\text{True Positives (correctly predicted positive cases)}$
$\mathit{TN}$ $:\text{True Negatives (correctly predicted negative cases)}$
$\mathit{FP}$ $:\text{False Positives (incorrectly predicted as positive)}$
$\mathit{FN}$ $:\text{False Negatives (incorrectly predicted as negative)}$
@@
To rigorously evaluate the performance of our proposed AugUNet1D architecture, we selected 16 diverse baseline models spanning multiple methodological paradigms and complexity levels. Our selection strategy was designed to provide comprehensive coverage of approaches used in EEG analysis and time series classification tasks. First, we included four traditional machine learning methods, namely Logistic Regression, K-Nearest Neighbors, Decision Tree, and Random Forest, which serve as foundational baselines and represent classical approaches to seizure detection that rely on learned feature representations from the input signals. Second, we incorporated several deep learning architectures of varying complexity: (1) standard CNN and PyramidalCNN for spatial feature extraction; (2) LSTM and bidirectional LSTM (biLSTM) for temporal modeling; and (3) ConvLSTM which combines both spatial and temporal processing. Third, we selected five specialized architectures that have demonstrated success in related physiological signal processing tasks: (1) EEGNet (specifically designed for EEG classification); (2) InceptionTime (a state-of-the-art time series classifier); (3) Xception (an efficient convolutional architecture); (4) vanilla U-Net (to isolate the contribution of our residual and augmentation enhancements); and (5) SalientSleepNet (a recent attention-based architecture for sleep stage classification). Finally, we included DETRtime, a transformer-based object detection model adapted for time series, to assess whether modern query-based detection paradigms are effective for dense temporal segmentation tasks. This diverse baseline selection enables the systematic evaluation of the relative contributions of different architectural components, including encoder-decoder structures, skip connections, residual learning, recurrent processing, and attention mechanisms to performance, and to determine which design principles are most critical for achieving robust cross-subject seizure detection performance.
@@
Performance scores (F1, recall, precision) of AugUNet1D and Twin Peaks were compared using paired samples t-tests. Differences in features (i.e., duration and peak frequency) among manually labeled SWDs, AugUNet1D-, and Twin Peaks-labeled events were tested using one-way repeated measures ANOVA with detection method as the main factor. Bonferroni post hoc tests were used to compare groups means among all possible combinations of the three detection methods. All statistical tests were performed in GraphPad Prism 8.0.1 (San Diego, CA). Data are reported as mean ± standard deviation (SD) unless otherwise stated.
@@
All EEG signals were identically preprocessed. First, raw signals were min-max scaled over their amplitude range to normalize signal magnitudes across subjects. Next, signals were resampled to 100Hz using torchaudio.transforms.Resample to standardize temporal resolution. Finally, the continuous recordings were segmented into 20-second epochs (2000 time points each) to create manageable training examples. The training data from the eight training mice was randomly partitioned into 95% for model training and 5% for validation, with the validation set used to monitor performance and select optimal model checkpoints during training. Batch size was 32.
The AugUNet1D model was trained using the Adam optimizer with an initial learning rate of $10^{-3}$ for a maximum of 50 epochs. We employed a CosineAnnealingLinearWarmup learning rate scheduler to facilitate stable training and optimal convergence, which began with a linear warmup phase over the first 500 steps to gradually increase the learning rate from zero to the initial value, preventing early training instability. Following warmup, the scheduler applied cosine annealing over 1000-step cycles with a minimum learning rate of $10^{-5}$ and a decay factor (gamma) of 0.9, allowing the learning rate to oscillate while gradually decreasing over successive cycles to enable fine-tuning of model parameters. To prevent overfitting and reduce unnecessary computational costs, we implemented early stopping with a patience of 10 epochs that monitored the validation Dice loss and terminated training if no improvement was observed for 10 consecutive epochs. This training configuration balanced exploration during early training phases with exploitation during later stages, whereas the early stopping criterion ensured that models were evaluated at their optimal generalization point rather than after potential overfitting. We report results as mean and standard deviation over three runs with random initialization.
@@ Results
Here, test the performance of a variety of traditional ML and neural network-based methods.
We specifically evaluated our proposed AugUNet1D model against 15 baseline methods and the Twin Peaks approach across 10 held-out test mice, using precision, recall, and F1-score as our primary evaluation metrics. The baseline methods spanned four methodological categories: (1) traditional machine learning algorithms (Logistic Regression, K-Nearest Neighbors, Decision Tree, Random Forest); (2) pure deep learning architectures (CNN, PyramidalCNN, LSTM, biLSTM, ConvLSTM); (3) specialized physiological signal processing networks (EEGNet, InceptionTime, Xception, U-Net, SalientSleepNet); and (4) transformer-based detection models (DETRtime). Additionally, we compared AugUNet1D against Twin Peaks, the previous state-of-the-art non-machine learning method on this dataset. Figure 5 shows the models’ overall average performance metrics (i.e. F1, recall, and precision scores). Table 2 presents the complete performance comparison across all methods and all 10 test mice.
Figure 5: Average performance metrics (F1, Recall, Precision) across all recordings in the test dataset. Left-to-right order of bars matches the top-to-bottom order in the legend. Performance by all methods is shown here, except for LSTM and biLSTM, because they yielded zero for all three metrics.
Table 2: Table shows the performance metrics for each model tested across all 10 of our test mice. The ‘avg’ columns contains the mean and standard deviation over all mice. We can see that our method performs the best on our dataset.
@@
Table 2 shows that our proposed AugUNet1D performed best, with an average F1-score of 0.90 ± 0.01, representing a 29% relative improvement over the Twin Peaks baseline (0.69 F1-score) and a 70% relative improvement over the best general deep learning baseline. Notably, AugUNet1D demonstrated exceptional cross-subject consistency with standard deviations of only 0.01-0.02 across all metrics, the lowest variance among all evaluated methods. The model achieved robust individual performance across all test subjects, with F1-scores ranging from 0.87 to 0.93, indicating successful generalization to held out mice. Furthermore, AugUNet1D maintained a balanced precision-recall trade-off with 0.91 precision and 0.90 recall, compared to Twin Peaks’ 0.67 precision and 0.81 recall, suggesting that our model achieves superior detection accuracy without sacrificing either sensitivity or specificity.
@@
The baseline methods exhibited dramatically different performance profiles that reveal important insights about architectural requirements for cross-subject seizure detection. Traditional machine learning methods performed poorly across the board, with F1-scores ranging from 0.16 for Logistic Regression to 0.33 for Decision Tree. Despite their consistent performance (low variance), these methods fundamentally struggled with the cross-subject temporal segmentation task likely due to their limited capacity to learn complex hierarchical representations from raw EEG signals. The failure of these methods underscores the necessity of deep learning approaches to solve this challenging problem.
@@
Surprisingly, pure recurrent architectures including LSTM, biLSTM, and ConvLSTM completely failed to learn effective representations, achieving 0.00 F1-scores across all test mice. This catastrophic failure suggests that recurrent networks, despite their theoretical suitability for sequential data, require careful initialization and architectural support to handle cross-subject generalization during high-frequency EEG signals. By contrast, standard convolutional architectures achieved moderate success, with vanilla CNN reaching a 0.47 ± 0.05 F1-score and PyramidalCNN slightly improving to an F1-score of 0.53 ± 0.04. However, these models exhibited high variance in precision (0.74 ± 0.21 for CNN), suggesting inconsistent performance across different subjects. The modest improvement of PyramidalCNN over vanilla CNN indicates that hierarchical feature extraction provides some benefit, but still falls far short for robust cross-subject detection.
@@
Among specialized architectures, vanilla U-Net achieved the highest F1-score of 0.59 ± 0.15, emerging as the strongest general deep learning baseline. However, its high standard deviation of 0.15 reveals significant instability across subjects, highlighting the importance of our residual connection modifications. InceptionTime and Xception both achieved identical performance, demonstrating that multi-scale temporal feature extraction and efficient convolutional designs alone are insufficient to overcome the cross-subject generalization challenge. Most surprisingly, EEGNet achieved only 0.19 ± 0.10 F1-score despite being specifically designed for EEG classification tasks. SalientSleepNet, despite incorporating attention mechanisms, achieved 0.47 ± 0.36 F1-score with extremely high variance (0.36 standard deviation), indicating that attention alone does not guarantee robust cross-subject generalization.
@@
Recently, "Twin Peaks" was developed for detecting SWDs generated by the WAG/Rij rat model of absence epilepsy, an approach that showed strong performance[23]. To determine how AugUNet1D compares to Twin Peaks at a granular resolution, we applied both methods to our test dataset and compared performance metrics as well as features of detected events (i.e. durations and peak frequencies).
@@
In these comparisons, performance was measured on a whole event basis. That is, if a detected event had any overlap with a manually labeled event, it was considered a true positive. We classified events with no such overlap at all as false positives (Figure 6A). Similarly, only manually labeled events that were missed entirely were considered false negatives (Figure 6B). Using this approach AugNet1D had a significantly higher mean F1 score (0.95 ± 0.04) than Twin Peaks (0.71 ± 0.15; t(9) = 5.34, p < 10-3). AugNet1D also showed higher precision (0.96 ± 0.03) than Twin Peaks (0.58 ± 0.18; t(9) = 7.38, p < 10-4). Twin Peaks exhibited a slightly higher recall (0.96 ± 0.07) than AugUNet1D (0.94 ± 0.06; t(9) = 3.95, p = 0.003), but the magnitude of the effects observed for F1 and precision scores were far larger. Because F1 scores balance precision and recall, we conclude that AugUNet1D outperformed Twin Peaks in our test dataset, having better F1 scores in every recording therein (Figure 6C).
@@
Figure 6: False positives, negatives, confusion matrices and F1 scores for AugUNet1D and Twin Peaks. A) Example of false positive in the dotted blue box. B) Example of false negative in the dotted gray box. Blue lines correspond to events detected by AugUNet1D. Gray lines correspond to manually labeled events in A and B. C) Confusion matrices and F1 scores for AugUNet1D and Twin Peaks prediction approaches for each mouse in the test dataset. Proportions in colored panels are as follows: Top left, true positives to all positives. Top right, false positives to all positives. Bottom left, false negatives to all negatives. Bottom right, true negatives to all negatives. These values were calculated on a per event basis, meaning that, for example, a predicted SWD was considered true if it had any overlap with an manually labeled SWD and false otherwise.
@@
We also systematically compared features of events detected by AugUNet1D and Twin Peaks to determine if either method has a bias towards particular event types. Specifically, we compared the peak frequency of detected events and the average duration of events, which, in our original training data, were similar across time and across mice. A one-way repeated measures ANOVA revealed a significant effect of detection method on event duration (F(2,18) = 92.7, p < 10-4; Figure 7A). Post hoc comparison showed that Twin Peaks detected events that were significantly briefer (2.35 ± 0.74s) than those detected either manually (4.88 ± 0.89s; t(9) = 11.7, p < 10-3) or by AugUNet1D (4.93 ± 0.82s; t(18) = 11.9, p < 10-3; Figure 7B). One-way repeated measures ANOVA also showed a significant effect of detection method on peak frequencies (F(2,18) = 18.25, p < 10-4; Figure 7C). Twin Peaks detected events with higher frequencies (6.58 ± 0.42Hz) than those detected by AugUNet1D (6.13 ± 0.14Hz; t(18) = 5.26, p < 10-3) or manually (6.14 ± 0.13Hz; t(9) = 5.2, p < 10-3; Figure 7D). There were no significant differences between AugUNet1D and manual detection in event duration (t(18) = 0.2, p > 0.99) or peak frequency (t(18) = 0.06, p > 0.99). To summarize, in addition to performing better in terms of overall F1 score, AugUNet1D also detected events with features similar to manually labeled SWDs. By contrast, Twin Peaks detected events with higher peak frequencies and shorter durations.
@@
Figure 7: Twin Peaks method shows a bias for shorter events with higher peak frequencies. A) Mean durations for events detected manually, with AugUNet1D, or with Twin Peaks in all ten mice in the test dataset. B) Bar heights represent mean durations across all mice. Open circles show the data for each mouse. C and D) Same as A and B but for peak frequencies of detected events. Error bars show ± standard deviation. *** p < 10-3
@@
We next asked whether events were erroneously detected during sleep or during electrographic noise induced by movement, poor mechanical connection, cable swinging, or during sleep. Briefly, noise epochs were defined by blocks of time filled with artificially high amplitude EEG signal (see Materials and Methods, Figure 8A). These recordings had very few noise epochs overall, making it difficult to estimate how these three methods performed during those periods. One-way repeated measures ANOVA revealed no significant effect of detection method on the proportion of events during noise (F(2,18) = 2.411, p = 0.214; Figure 8B).
Sleep epochs were identified by an increased amplitude in the 0.1 to 4Hz-bandpass filtered EEG signal (see Materials and Methods; Figure 8C). One-way repeated measures ANOVA revealed a significant effect of detection method on the proportion of events during sleep (F(2,18) = 35.06, p < 10-4). Post hoc comparisons showed that a greater proportion of events during sleep were detected by Twin Peaks (0.46 ± 0.15) than by manual labeling (0.17 ± 0.07; t18 = 7.36 p < 10-4) and AugUNet1D (0.18 ± 0.017; t18 = 7.14 p < 10-4). There was no difference between manually detected and AugUNet1D-detected events (t(18) = 0.22, p > 0.99, Figure 8D). In conclusion, short, non-SWD events are erroneously detected by Twin Peaks during sleep.
Figure 8: Cross-method comparison of detected events during noise and sleep. A) Example of the Twin Peaks method falsely detecting two events during periods with high-amplitude cable artifact highlighted in yellow shading. B) Overall, very few events were detected during noise epochs, but Twin Peaks detected more than AugUNet1D and manual labeling. B) The proportion of events that Twin Peaks detects during sleep is higher than the number of manually detected events and those detected by AugUNet1D. B) Example trace with Twin Peaks falsely detecting many events during sleep which is indicated in blue shading D) Twin Peaks detects many more events during sleep than manual labeling and AugUNet1D. **** p < 10-4
@@
AugUNet1D uses several data augmentation strategies during training to diversify the dataset and improve generalization to new, unseen data. We conducted an ablation study to assess the individual and combined contributions of each data augmentation strategy: Gaussian noise injection, signal inversion, and amplitude scaling—on the cross-subject seizure detection performance of our residual 1D U-Net architecture. The results reveal critical insights into the role of augmentation in achieving robust generalization. Results are presented in Table 3.
Table 3: Model performance for different chosen augmentations. We can see scaling provides the most benefit for our use case.
@@
Without any data augmentation, the model achieved high precision (0.9464) but catastrophically low recall (0.2755), resulting in a poor F1-score of 0.4268. This severe imbalance indicates that the unaugmented model learned to be extremely conservative in its predictions, likely defaulting to predicting the majority (non-seizure) class to minimize false positives. Although the model could identify seizures with high confidence when it made positive predictions, it missed the vast majority of bona fide seizure events, rendering it clinically ineffective.
Including Gaussian noise augmentation alone into the model provided only marginal improvement, increasing recall to 0.3116 and the F1-score to 0.4698 while maintaining high precision (0.9544). This modest benefit suggests that noise injection helps the model become slightly more robust to signal variations, but does not fundamentally address the class imbalance problem. Signal inversion augmentation showed slightly greater impact, improving recall to 0.3631 and F1-score to 0.5267 with precision of 0.9591.
By contrast, amplitude scaling augmentation alone produced a dramatic transformation in model behavior, achieving a recall of 0.8442 and F1-score of 0.8609. While precision decreased moderately to 0.8783, the substantial gain in recall demonstrates that scaling augmentation is by far the most critical component for cross-subject generalization.
Ultimately, combing all three augmentation strategies yielded the best overall performance with an F1-score of 0.8848, achieving an optimal balance between precision (0.9092) and recall (0.8618). The combined approach outperformed even amplitude scaling alone, indicating that while scaling addresses the primary challenge of amplitude variability, the additional regularization resulting from noise injection and the polarity invariance from signal inversion provide complementary benefits that further improve model robustness.